# MnBr2 on the graphene on Ir(110) substrate: growth, structure, and super-moiré

### v1

@@ II.2 Morphological Evolution with Temperature in Growth and Annealing
Figure 2 compares the morphology of MnBr2 as a function of growth temperature. Figure 2(a), taken at 77 K after growth at 80 K with unavoidable annealing to $\approx 150$ K in between due to sample transfer, reveals fractal island growth with smooth edges and no preferential orientation. The smooth edges suggest limited edge diffusion during the annealing period. In the absence of annealing at $\approx$ 150 K, the islands would probably have retained even more irregular fractal shape. As shown in the inset of Figure 2(a), most of the islands exhibit a single well-defined moiré orientation, except when grown together; an example of this is found in the upper left corner of the contrast-enhanced inset. The inset shows stripes of the Gr/Ir(110) substrate moiré. The average periodicity is 3.3 nm [56], although due to the magnification effect of the Gr/Ir(110) moiré for strain and defects, substantial variations occur around this average. In any case, even when the MnBr2 moiré stripes are parallel to those of the substrate, their wavelength is substantially smaller. Second-layer islands are largely absent. The average distance between islands of about 25 nm is much smaller than the typical separation of about 150 nm between steps or other surface defects. This implies the dominance of homogeneous nucleation.
@@
Figure 2(b), acquired at 300 K after growth at 300 K, shows a segment of a large dendritic island. While the main visible branch of the dendrite grows from bottom to top [direction labeled 1 in the inserted sketch in Figure 2(b)], secondary branches grow away from it [directions labeled 2], maintaining angles of approximately 120. These angles correspond to the preferred growth directions, which are consistent with the three-fold symmetry of the crystal structure. The island edges are smooth on a much larger length scale than in Figure 2(a) and display undulations. Sharp grooves were observed between the secondary branches. Their origin can be traced back to the limited molecule supply from the diffusion field, which is preempted by the rapidly growing secondary branches and islands in the second layer. Some of the grooves are pinched, causing holes. The onset of tertiary branches is visible in Figure 2(b) [directions labeled 3]. From the dendritic appearance, the rather smooth edges, and the simultaneous existence of grooves, it must be concluded that step edge diffusion is effective on length scales of the order 5 nm, but limited on large length scales of the order 50 nm. Searching the sample by moving the STM scan frame shows dendritic growth all over, but the substrate steps and defects give rise to heterogeneous nucleation, making a quantitative assessment of the nucleation density impossible to achieve. Qualitatively, the strong reduction in the island number density with increasing growth temperature is consistent with classical nucleation theory [59]. A large number of second-layer islands nucleated on top of the dendritic island. This observation implies that the diffusion of MnBr2 molecules on the first-layer MnBr2 islands is hampered compared to diffusion on Gr/Ir(110), likely due to a larger activation barrier for molecule migration. Second-layer islands are seen to be reduced in number density next to the steps of the first-layer islands. This implies that for molecules diffusing on a first-layer island, the edge acts as a sink for molecules by incorporating them. Consequently, if there is any Ehrlich-Schwoebel barrier [60, 61, 59] for the descent of molecules from the first-layer island, it must be very small.
@@
Figure 2(c) taken at 300 K and grown at 400 K shows two first-layer islands grown together. The right island is a good example of dendritic-skeletal growth [59]. Its envelope is triangular, but its triangular sides are concave. The diffusion field has the strongest concentration gradient at the triangle tips; thus, the tips receive the largest supply of molecules. Although diffusion is efficient, transport along the edge is not fast enough to keep the edges straight, causing concave edges. Due to the higher growth temperature, the nucleation density of second-layer MnBr2 islands is smaller compared to the 300 K case. The zone denuded of second-layer islands adjacent to the first-layer island edges has grown in width, consistent with enhanced molecule diffusion at higher temperature. The second-layer islands are compact, and if not coalesced, of hexagonal to triangular in shape, since, due to their small size, diffusion is efficient in keeping the edges straight.
@@
Taken together, the inferred picture of island nucleation and shape evolution is consistent with a strong increase with the growth temperature of (i) molecule diffusion on Gr/Ir(110) and (ii) step edge diffusion. The much higher diffusion coefficient of molecules on Gr/Ir(110) compared to that on the first-layer MnBr2 islands implies that homogeneous nucleation has already ceased for the first-layer islands at 300 K, while for the second-layer islands, it remains homogeneous throughout the investigated temperature range. The higher molecular diffusion coefficient on Gr/Ir(110) also indicates a weaker interaction between MnBr2 and Gr than between layers of MnBr2.
@@
Figure 3: (a)-(d) The MnBr2 sample of Figure 2(b) grown at 300 K is annealed in isochronal steps of 120 s successively to 370 K, 470K̇, 570 K, and 670 K, respectively. The white arrows in (c) highlight the two pinning centers. Imaging is conducted 300 K with $V_{\mathrm{b}}$ = -2 V, $I_{\mathrm{t}}=50$ pA, Images sizes: 160 nm $\times$ 160 nm.
@@
II.3 Super-moiré
Figure 4: Ball model of super-moiré formation through overlay of the Ir(110) surface lattice (starting left), the Gr lattice (starting right), and one Br sublattice of MnBr2 (starting at the bottom). The MnBr2 lattices are rotated counterclockwise in (a) by $0.5^{\circ}$ and in (b) by $6.5^{\circ}$ with respect to the Ir[11¯0][$\bar{1}$0] direction.
The first insight into the moiré pattern in the MnBr2/Gr/Ir(110) system is obtained by overlaying ball-model representations of the three constituent lattices, as shown in Figure 4. The Ir(110) surface lattice (left) overlaps in the central area of both subfigures with the Gr honeycomb lattice (right). On these lattices, one Br (or equivalently the Mn) sublattice is overlaid (green balls, lattice shifted a bit down) at two different angles. On the left of each subfigure, a clean stripe moiré pattern is observed, and on the right, a hexagonal moiré pattern evolves, as indicated by an orange double arrow and a pink rhombus, respectively. In the region where all three lattices overlap, the stripe pattern survives but becomes modulated along the stripes due to the additional MnBr2 lattice. The stripe pattern is the dominant feature in the moiré patterns highlighted in Figure 1(b) and 2(a). The orientation and width of the stripes in the ball model are well matched to those highlighted in Figure 1(b). The fit parameter used was the slight misorientation of the MnBr2 lattice with respect to Gr/Ir[110], with angles of $0.5^{\circ}$ and $6.5^{\circ}$.
Figure 5: Fast Fourier transformation (FFT) analysis of the MnBr2/Gr/Ir(110) super-moiré. (a) Moiré-resolution STM topography of MnBr2. (b) In the upper left, the FFT of the STM image shown in (a) is presented, to the right of which is an enlarged, color-coded schematic model of the FTT. The relevant $\vec{k}$ vectors are indicated. (c-e), (g), and (h) are inverse FFTs using the spots represented by (c) pink, (d) orange, (e) blue, (g) purple, and (h) green dots in (b). (f) Sum of images shown (d) and (e). STM imaging parameters in (a) are $V_{\mathrm{b}}=-1$ V and $I_{\mathrm{t}}=50$ pA. Image size is 50 nm $\times$ 50 nm.
The super-moiré patterns in the MnBr2/Gr/Ir(110) system exhibit additional complex features beyond those visualized in the ball model. To identify the components of the super-moiré, we performed a fast Fourier transform (FFT) analysis of the high-resolution topography of the super-moiré, as shown in Figure 5(a), which is taken on the island visible in the upper part of Figure 1(b). The height profile reveals a substantial corrugation of $\approx$ 0.7 Å (see Figure S2 of SI). The FFT of Figure 5(a), presented in Figure 5(b), displays multiple peaks. The central part of the FFT limited to small $\vec{k}$ or large wavelengths, is represented in the upper left corner of Figure 5(b). A color-coded, enlarged, and schematic version of the pattern is shown in Figure 5(b). The inverse FFT analysis of sets of equally colored peaks provides insight into the main components of the super-moiré. The inverse FFT is shown in Figure 5(c) when selecting the hexagonal FFT peaks (encircled in pink). The image represents the hexagonal moiré between Gr and MnBr2 already visible in the ball model of Figure 4(a) by the pink rhombus. It has a lattice parameter of 2.5 nm. The inverse FFT is shown in Figure 5(d) when the two most pronounced Fourier peaks (orange) are selected. The image represents the interference between MnBr2 and Ir(110). It has a row spacing of 2.7 nm and gives rise to a pronounced line pattern in the MnBr2/Gr/Ir(110) super-moiré. This line pattern is clearly visible in Figure 4(a) on the left, where the Ir(110) surface rows overlap with the MnBr2 lattice. The strength of this pattern in the super-moiré is remarkable considering that Ir(110) and MnBr2 are separated by the Gr layer. The origin of this phenomenon is further analyzed in this section. Selecting the blue Fourier peaks is shown in Figure 5(e). The stripes have the larger periodicity of the Gr/Ir(110) moiré, which is 3.3 nm [56]. This periodicity is also visible in between the MnBr2 islands in the contrast-enhanced inset of Figure 2(a). The superposition of the two stripe patterns results in a beating pattern with a period of 13.5 nm. Superposition can be achieved by simply adding the two images, as shown in Figure 5(f), or by selecting the green spots in the FFT to create an inverse FFT, as shown in Figure 5(g). The green spots are due to the difference in the wave vector of the blue and orange spots [or the sum of the blue and orange $\vec{k}$ vectors in Figure 5(b)]. The modulation with a period of 13.5 nm is visible in the STM topograph of Figure 5(a). Finally, when selecting the purple spots resulting from the addition of $\vec{k}[\mathrm{Gr/Ir(110}]+\vec{k}[\mathrm{MnBr_{2}/Ir(110}]$ and similar sums (or differences), the inverse FFT is shown in Figure 5(h). The beating amplitudes of both, the vertical and stripe patterns can be recognized. The horizontal stripes are phase-shifted by $\pi$ between the segments of 13.5 nm periodicity, as in the STM topograph of Figure 5(a).
Figure 6: (a) Atomically resolved moiré of Gr/Ir(110). The inset shows the height profile along the red line. (b) DFT-calculated charge density isosurface with $2.1\times 10^{-10}$ electrons/Å3 in an average height of $3.6$ Å above the surface. The grayscale encodes the variation in isosurface height. (c) Ball model of atomic arrangement in the supercell. Navy blue: bottom-layer Ir atoms; light blue balls: top-layer Ir atoms; gray balls: C atoms. A black rhombus in all subfigures indicates the supercell size in the DFT calculations. STM imaging parameters in (a) $V_{\mathrm{b}}=100$ mV and $I_{\mathrm{t}}=1$ nA. Image size is 10 nm $\times$ 8 nm.
After establishing the main elements of the super-moiré, it remains to be understood why the Fourier component corresponding to a "virtual" moiré of MnBr2 and Ir(110) is so pronounced in the super-moiré, although the two lattices are not in contact. The answer is encoded in the atomically resolved STM topography of Gr on Ir(110), as shown in Figure 6(a). The topograph clearly shows the rectangular $3.3\,\mathrm{nm}\times 1\,\mathrm{nm}$ unit cell of the Gr/Ir(110) moiré and the corresponding stripe periodicity. However, in addition to this, the Ir(110) atomic row spacing of 3.84 Å is imprinted onto the moiré. The densely packed Ir(110) rows run vertically under Gr in Figure 6(a). The imprinted periodicity is evident from the line profile in the inset of Figure 6(a), which is on a horizontal ridge along the red line in Figure 6(a) and displays a 3.84 Å periodicity. The same Ir row periodicity is also observed in the horizontal grooves. Our findings are supported by the DFT calculations of the total charge density of the Gr/Ir(110) moiré displayed in Figure 6(b). The total charge density was calculated using DFT for the supercell indicated by a black rhombus in Figure 6(b), of which the atomic structure is shown in Figure 6(c). The charge density isosurface with $2.1\times 10^{-10}$ electrons/Å3 is at an average height of 3.6 Å above the surface. Darker gray represents smaller and lighter gray larger heights. The total charge density isosurface not only resembles the STM topograph quite well but also has the Ir atomic row periodicity imprinted as strongly as in the STM topograph. In particular, the Gr lattice is visually absent in the STM topography in Figure 6(a) and the DFT-calculated charge density in Figure 6(b). In the next section, we will see that MnBr2 is a large-band-gap insulator, for which our tunneling conditions are within the bandgap. Thus, the STM signal is dominated by tunneling into the interface of the bottom Br sublattice of MnBr2 in contact and hybridized with the electronic structure of Gr/Ir(110) shown in Figure 6. In this sense, the super-moiré observed on the MnBr2 islands is an electronic super-moiré structure situated below the MnBr2 surface. The dominance of the Fourier components related to $\vec{k}[\mathrm{MnBr_{2}/Ir(110)}]$ is a consequence of the periodicity of the top layer Ir(110) rows, which are present with a large modulation in the Gr/Ir(110) moiré.
II.4 Apparent Height and Density of States
@@
The apparent STM height of single-layer MnBr2 on Gr/Ir(110) is plotted in Figure 7(a) as a function of the bias voltage $V_{\mathrm{b}}$. Over a large range from -4 V to 2 V it is rather insensitive to $V_{\mathrm{b}}$ with an almost constant value of $\sim$4 Å. Only above $V_{\mathrm{b}}=2\,V$ it increases, reaches at $V_{\mathrm{b}}$ = 3.2 V a maximum of 5.95 $\pm$ 0.05 Å and then decreases again. We note that the precise apparent heights measured depend on $I_{\mathrm{t}}$ and the tip state, but its $V_{\mathrm{b}}$-dependence, as represented in Figure 7(a) is generic. The measured height did not reach the experimental c-axis lattice parameter of bulk MnBr2, which is 6.271 Å[57]. The situation is similar for second-layer islands. The apparent heights are in the range of 4 Å- 6 Å, dependent on $V_{\mathrm{b}}$, $I_{\mathrm{t}}$, and the tip state. For an example height profile, see Figure S1 of the SI. The situation is similar to that of CoCl2, where different apparent heights ranging from 4.8 Å to 8.0 Å and a strong bias dependence have been reported [11, 10, 15].
@@
Figure 7(b) displays a differential conductance ($\mathrm{d}I/\mathrm{d}V$) spectrum acquired from a defect-free region more than 10 nm from the edges of single-layer MnBr2 (blue data). The differential conductance is, in the first approximation, proportional to the local density of states of the sample. The conduction band edge (CBE) is located at $V_{\mathrm{b}}=(-2.5\pm 0.1)$ V and the valence band edge (VBE) at $V_{\mathrm{b}}=(1.9\pm 0.1)$ V. The resulting total gap of $(4.4\pm 0.2)$ eV makes MnBr2 an insulator, consistent with previous DFT calculations that found a band gap of 4.21 eV [24] or 3.95 eV [25].
Figure 7(b) also displays the $\mathrm{d}I/\mathrm{d}V$ spectrum of bilayer MnBr2 (red data). The CBE is located at $V_{\mathrm{b}}=(-3.0\pm 0.1)$ V and the VBE at $V_{\mathrm{b}}=(2.1\pm 0.1)$ V. The total resulting gap of $(5.1\pm 0.2)$ eV is larger than the band gap of the single-layer, contrary to the case of MoS2 [62]. The larger band gap of the bilayer is at variance with DFT calculations. They yielded a smaller band gap for the MnBr2 bulk than for the single-layer [25]. However, a similar trend has been reported for CoCl2 grown on HOPG (see Supporting Information of ref. 11), where the bilayer displays a larger band gap than the single-layer.
Evidently, the strong variation in the apparent height, as shown in Figure 7(a), is related to the electronic structure of MnBr2. In the range of the band gap between $V_{\mathrm{b}}=(-2.5\pm 0.1)$ V and $V_{\mathrm{b}}=(1.9\pm 0.1)$ V tunneling is between the tip and the corrugated moiré substrate Gr/Ir(110). The latter may be hybridized with the Br sublattice in contact with it, thus providing a mechanism for the MnBr2 layer to enter into a lateral modulation of the tunneling probability. The tunneling barrier within MnBr2 is defined by its CBE being lower than that of the vacuum barrier. This gives rise to the apparent non-zero height of MnBr2. The modulation of the barrier height by the distribution of electronic states in the CBE is consequently the second mechanism by which MnBr2 modulates the tunneling probability laterally. Both mechanisms are primarily electronic in origin and are not related to the variation of the height of the nuclei in the MnBr2 layer with lateral position. This is supported by the fact that the apparent corrugation measured in the supermoiré with 0.7 Å  is much larger than the corrugation of the Gr/Ir(110) moiré with 0.2 Å [see Figure 6(c)].

### v2

@@ II.2 Morphological Evolution with Temperature in Growth and Annealing
Figure 2 compares the morphology of MnBr2 as a function of growth temperature. Figure 2(a), taken at 77 K after growth at 80 K with unavoidable annealing to $\approx 150$ K in between due to sample transfer, reveals fractal island growth with smooth edges and no preferential orientation. The smooth edges suggest limited edge diffusion during the annealing period. In the absence of annealing at $\approx$ 150 K, the islands would probably have retained an even more irregular fractal shape. As shown in the inset of Figure 2(a), most of the islands exhibit a single well-defined moiré orientation, except when grown together; an example of this is found in the upper left corner of the contrast-enhanced inset. One island in the inset does not display a visible moiré, for which the reason will be discussed in Section Super-moiré. The inset also shows the moiré of the Gr/Ir(110) substrate, which is always present, but much better visible for the specific bias used to take Figure 2(a). The average moiré stripe periodicity is 3.3 nm [Kraus22], although due to the magnification effect of the Gr/Ir(110) moiré for strain and defects, substantial variations occur around this average. In any case, even when the MnBr2 moiré stripes are parallel to those of the substrate, their wavelength is substantially smaller. Second-layer islands are largely absent. The average distance between islands of about 25 nm is much smaller than the typical separation between steps or other surface defects of about 150 nm. This implies the dominance of homogeneous nucleation.
@@
Figure 2(b), acquired at 300 K after growth at 300 K, shows a segment of a large dendritic island. While the main visible branch of the dendrite grows from bottom to top [direction labeled 1 in the inserted sketch in Figure 2(b)], secondary branches grow away from it [directions labeled 2], maintaining angles of approximately 120. These angles correspond to the preferred growth directions, which are consistent with the three-fold symmetry of the crystal structure. The island edges are smooth on a much larger length scale than in Figure 2(a) and display undulations. Sharp grooves were observed between the secondary branches. Their origin can be traced back to the limited molecule supply from the diffusion field, which is preempted by the rapidly growing secondary branches and islands in the second layer. Some of the grooves are pinched, causing holes. The onset of tertiary branches is visible in Figure 2(b) [directions marked 3]. From the dendritic appearance, the rather smooth edges and the simultaneous existence of grooves, it must be concluded that step edge diffusion is effective on length scales of the order 5 nm, but limited on large length scales of the order 50 nm. Searching the sample by moving the STM scan frame shows dendritic growth all over, but the substrate steps and defects give rise to heterogeneous nucleation, making a quantitative assessment of the nucleation density impossible to achieve. Qualitatively, the strong reduction in the island number density with increasing growth temperature is consistent with classical nucleation theory [Michely04]. A large number of second-layer islands nucleated on top of the dendritic island. This observation implies that the diffusion of MnBr2 molecules on the first-layer MnBr2 islands is hampered compared to diffusion on Gr/Ir(110), likely due to a larger activation barrier for molecule migration. Second-layer islands are seen to be reduced in number density next to the steps of the first-layer islands. This implies that for molecules diffusing on a first-layer island, the edge acts as a sink for molecules by incorporating them. Consequently, if there is any Ehrlich-Schwoebel barrier [Ehrlich66, Schwoebel66, Michely04] for the descent of molecules from the first-layer island, it must be very small.
@@
Figure 2(c) taken at 300 K and grown at 400 K shows two first-layer islands grown together. The right island is a good example of dendritic-skeletal growth [Michely04]. Its envelope is triangular, but its triangular sides are concave. The diffusion field has the strongest concentration gradient at the triangle tips; thus, the tips receive the largest supply of molecules. Although diffusion is efficient, transport along the edge is not fast enough to keep the edges straight, causing concave edges. Due to the higher growth temperature, the nucleation density of second-layer MnBr2 islands is smaller compared to the 300 K case. The zone denuded of second-layer islands adjacent to the first-layer island edges has grown in width, consistent with enhanced molecule diffusion at higher temperatures. The second-layer islands are compact and, if not coalesced, of hexagonal to triangular shape, since, due to their small size, diffusion is efficient in keeping the edges straight.
@@
Taken together, the inferred picture of island nucleation and shape evolution is consistent with a strong increase with the growth temperature of (i) molecule diffusion on Gr/Ir(110) and (ii) step edge diffusion. The much higher diffusion coefficient of molecules on Gr/Ir(110) compared to that on the first-layer MnBr2 islands implies that homogeneous nucleation has already ceased for the first-layer islands at 300 K, while for the second-layer islands, it remains homogeneous throughout the investigated temperature range.
Figure S2 of the SI for 400 K growth of MnBr2 on Gr/Ir(111) shows qualitatively the same behavior as for the 400 K growth on Gr/Ir(110). Nucleation, second-layer nucleation, diffusion, and island shapes of MnBr2 appear to be similar on both substrates.
@@
Figure 3: (a)-(d) The MnBr2 sample of Figure 2(b) grown at 300 K is annealed in isochronal steps of 120 s successively to 370 K, 470K̇, 570 K, and 670 K, respectively. White arrows in (c) highlight two pinning centers. Imaging is conducted 300 K with $V_{\mathrm{b}}$ = -2 V, $I_{\mathrm{t}}=50$ pA, Images sizes: 160 nm $\times$ 160 nm.
@@
II.3 Apparent Height and Density of States
@@
The apparent STM height of single-layer MnBr2 on Gr/Ir(110) is plotted in Figure 4(a) as a function of the bias voltage $V_{\mathrm{b}}$. Over a large range from -4 V to 2 V it is rather insensitive to $V_{\mathrm{b}}$ with an almost constant value of $\sim$4 Å. Only above $V_{\mathrm{b}}=2\,V$ it increases, reaches at $V_{\mathrm{b}}$ = 3.2 V a maximum of 5.95 $\pm$ 0.05 Å and then decreases again. The maximal measured height at $V_{\mathrm{b}}$ = 3.2 V almost reaches the experimental c-axis lattice parameter of bulk MnBr2, which is 6.272 Å [Wollan58]. We note that the precise apparent heights measured depend on $I_{\mathrm{t}}$ and the tip state, but its $V_{\mathrm{b}}$-dependence, as represented in Figure 4(a) is generic.
Also for second-layer islands the apparent heights are in the range of 4 Å - 6 Å, dependent on $V_{\mathrm{b}}$, $I_{\mathrm{t}}$, and the tip state. For an example height profile of a second-layer island, see Figure S1 of the SI. The situation is similar to that of CoCl2, where different apparent STM heights ranging from 4.8 Å to 8.0 Å and strong bias dependencies were reported [liu2023, cai2023, Hao24]. Atomic force microscopy provides a height of 6.3 Å for CoCl2, reasonably close to the expected bulk value of 5.9 Å [Kerschbaumer25]. As discussed in the following, electronic effects are responsible for the variation in the apparent height measured by STM.
@@
Figure 4(b) displays a differential conductance ($\mathrm{d}I/\mathrm{d}V$) spectrum acquired from a defect-free region more than 10 nm from the edges of single-layer MnBr2 (blue curve). The differential conductance is, in the first approximation, proportional to the local density of states of the sample. The valence band edge (VBE) is located at $V_{\mathrm{b}}=(-2.5\pm 0.1)$ V and the conduction band edge (CBE) at $V_{\mathrm{b}}=1.9\pm 0.1$ V. The resulting total gap of $(4.4\pm 0.2)$ eV makes MnBr2 an insulator.
Screening through the Gr substrate may be expected to have renormalized our measured band gap to values smaller than of freestanding MnBr2, as was observed for instance for MoSe2 [Ugeda14] or MoS2 [vanEfferen22]. Consistent with this speculation, we observe a larger band gap for bilayer MnBr2, for which screening through the Gr substrate is reduced. As shown in Figure 4(b), for bilayer MnBr2 (red data) the VBE is located at $V_{\mathrm{b}}=(-3.0\pm 0.1)$ V and the CBE at $V_{\mathrm{b}}=(2.1\pm 0.1)$ V. The total resulting gap of $(5.1\pm 0.2)$ eV is significantly larger than the band gap of the single-layer. A similar trend has been reported for CoCl2 grown on graphite (see the Supporting Information of ref. liu2023), where the bilayer displays a larger band gap than the single-layer. For freestanding single-layer MnBr2 previous DFT calculations found a band gap of 4.21 eV [Luo20] or 3.95 eV [Jesus25] and an even smaller band gap for MnBr2 bulk. These results are consistent with MnBr2 being an insulator, but seem to underestimate the band gap.
Evidently, the strong variation in apparent height, as shown in Figure 4(a), is related to the electronic structure of MnBr2. In the range of the band gap between $V_{\mathrm{b}}=(-2.5\pm 0.1)$ V and $V_{\mathrm{b}}=(1.9\pm 0.1)$ V tunneling is between the tip and the moiré substrate Gr/Ir(110). The tunnel barrier is composed of the vacuum barrier between the tip and MnBr2 and the barrier within MnBr2 between its two surfaces facing vacuum and Gr/Ir(110). The barrier within MnBr2 is defined by its CBE. Since the CBE is lower in energy than the vacuum level, the presence of MnBr2 enhances the tunneling probability and thus gives rise to a non-zero height of MnBr2 in the constant current mode.
When $V_{\mathrm{b}}$ decreases below 2.5-2.5 V, a new tunneling channel opens: electrons can then tunnel also out of the MnBr2 valence band. However, it is far below the Fermi level of the sample. Compared to electrons close to the sample Fermi level, their effective tunneling probability is lower by far, because of a much larger effective barrier height. Thus, the effect of tunneling out of the deep valence band states 2.5-2.5 eV below the Fermi level is neglectable compared to tunneling out of Gr/Ir(110) states near the Fermi level. Therefore, no significant change in the apparent barrier height is observed.